# MultiFinBen: A Multilingual, Multimodal, and Difficulty-Aware Benchmark for Financial LLM Evaluation

### v2

@@ 4 Results
(c) Performance across difficulty levels: Easy, Medium, Hard.
@@
Monolingual models underperform in multilingual settings.
English-only models, such as finma-7b-full (8.91%, 15nd), consistently trail bilingual/multilingual models like Llama3.1-XuanYuan (10.42%, 12th), FinMA-ES (9.65%, 14th), and plutus (11.82%, 10th). This performance gap widens between FinMA-7b-full and plutus further in bilingual (69.24% vs. 91.59%) and multilingual (3.10% vs. 7.24%) evaluation scenarios, reflecting the critical impact of cross-lingual capacity. Moreover, language-specific performance aligns closely with training data coverage. High-resource languages (e.g., English, Chinese, Japanese, Spanish) are dominated by generalist frontier models such as GPT-4o and Llama-4, while domain-specialized models like plutus achieve superior results in low-resource languages, as evidenced in Greek (60.19% vs. Llama-4’s 48.95%). These findings establish that multilinguality is not merely an auxiliary feature but a foundational requirement for LLMs in finance, positioning MultiFinBen as the first comprehensive benchmark to diagnose and drive progress in multilingual financial language understanding.
Structured difficulty-aware benchmarking facilitates dynamic evaluation.
While early models such as finma-7b-full perform competitively on easy tasks (49.48%), performance declines sharply on medium (22.01%) and hard tasks (9.49%), with newer models like GPT-4o showing improvements yet continuing to struggle. To address the lack of coverage in existing benchmarks, we introduce PolyFiQA-Easy and PolyFiQA-Expert, which are the first multilingual financial datasets, and the first financial OCR task, explicitly designed to target underrepresented linguistic and modality challenges. These datasets, initially created to fill specific gaps, emerged as some of the most difficult evaluation tasks, with Deepseek-V3 achieving only 42.58% and 31.40% on PolyFiQA-Easy and PolyFiQA-Expert, and overall model averages of 7.50% and 5.61%, respectively. This finding highlights how MultiFinBen’s structured design not only reveals current model limitations but also serves as a practical framework to systematically guide the development of new datasets and scenarios, ensuring benchmarks evolve in step with model capabilities and real-world financial demands.
@@ 5 Conclusion
We introduce MultiFinBen, the first multilingual and multimodal benchmark designed to systematically evaluate real-world financial application under difficulty-aware settings. Our analysis reveals persistent performance gaps across modalities, languages, and difficulty levels, with even GPT-4o showing substantial limitations, particularly in complex multilingual and multimodal scenarios. Through our modality-language-task framework and the introduction of novel datasets, including PolyFiQA-Easy, PolyFiQA-Expert, EnglishOCR, and SpanishOCR, we stress-test existing models and expose nuanced limitations in long-form reasoning, audio understanding, OCR, and multilingual comprehension. The sharp performance drop on multilingual tasks demonstrates that scaling alone is insufficient for global financial applications. MultiFinBen provides a comprehensive, dynamic evaluation platform and offers a foundation for future research toward inclusive, multilingual, and multimodal financial AI. By releasing all datasets and evaluation scripts, we aim to foster reproducible and impactful research and encourage the community to develop the next generation of financial LLMs capable of robust, cross-modal, and cross-lingual reasoning in real-world economic and policy contexts.
Acknowledgments and Disclosure of Funding
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. Shengyuan Colin Lin, Keyi Wang, and Xiao-Yang Liu Yanglet acknowledge the support from Columbia’s SIRS and STAR Program, The Tang Family Fund for Research Innovations in FinTech, Engineering, and Business Operations. Shengyuan Colin Lin and Xiao-Yang Liu Yanglet acknowledge the support from NSF IUCRC CRAFT center research grant (CRAFT Grant 22017) for this research. The opinions expressed in this publication do not necessarily represent the views of NSF IUCRC CRAFT.
Appendix A Related Work
A.1 Trend of Benchmark Development
With the rapid development of large language models, many traditional benchmark datasets have become insufficiently challenging to effectively distinguish between models of varying capacities [28]. Even relatively small-scale models now achieve comparable performance to much larger ones on these benchmarks, despite notable disparities in their real-world capabilities [22]. The FinBen benchmark [77] also reveal that certain datasets are unable to clearly distinguish differences in model capabilities by conducting thorough experiments and evaluates LLMs across diverse financial tasks.
This performance saturation has motivated increasing efforts to design more difficult and nuanced evaluation tasks across various domains [19, 73, 74]. In the general domain, the BIG-Bench dataset [67] compiles challenging tasks to stress-test LLMs. Subsequent work examines whether prompting techniques can enhance performance in such tasks. In the medical domain, Chen et al. [6] developed QA tasks requiring clinical reasoning and explanation, providing a more rigorous evaluation of LLMs’ capabilities in healthcare. In the financial domain, Zhao et al. [90] introduced datasets of varying difficulty. For simpler tasks, small models perform comparably to larger ones. However, for complex benchmarks like Complong, involving multi-table reasoning and long document processing, smaller models struggle while larger models maintain a lead. This highlights the need for more discriminative benchmarks to expose nuanced differences in model capabilities.
Building on these observations and the proliferation of financial datasets, some recent work has proposed criteria to filter and prioritize representative datasets for benchmarking [59, 3]. These efforts assess benchmarks based on implementation, design, documentation, maintenance, and construction to determine their effectiveness. However, these approaches do not systematically evaluate the impact of dataset difficulty or establish objective criteria for filtering datasets based on complexity.
A.2 Multilingual Textual Financial Benchmarks
Despite the growing number of financial benchmarks developed to assess the capabilities of large language models in domain-specific tasks, most existing benchmarks remain monolingual, predominantly focusing on English or Chinese. Notable examples include PIXIU [79], CFBenchmark [35], FinanceBench [27], and BizBench [32], each offering tasks such as question answering, text classification, and numerical reasoning, all rooted in financial documents or news within a single language setting. Some efforts have extended financial benchmarking to additional languages. For instance, [54] and [21] introduced benchmarks that encompass Greek and Japanese, broadening the linguistic focus beyond English and Chinese. Although works like [77] and [87] expand beyond English by including tasks in both English and Spanish, they treat the two languages as separate tasks rather than integrating them into a unified multilingual evaluation. This approach limits the ability to assess cross-lingual generalization within financial contexts.
Some recent studies have attempted to expand financial benchmarks by incorporating bilingual tasks. For example, DolFin [46] includes English-Chinese financial QA and reasoning tasks. However, its focus remains limited to translation-based tasks, which do not fully reflect real-world multilingual applications.
Currently, there is a lack of comprehensive financial benchmarks that simultaneously encompass monolingual, bilingual, and multilingual settings. Most existing resources either focus solely on monolingual tasks or limited bilingual tasks, without integrating them into a unified multilingual framework. This gap highlights the need for more inclusive and versatile evaluation resources that can assess model performance across diverse linguistic settings. To this end, our work introduce Our work addresses this limitation by proposing a benchmark that not only evaluates LLMs across a wider range of financial tasks, but also considers dataset value filtering and cross-lingual robustness, setting the stage for more comprehensive and discriminative benchmarking in the financial domain.
A.3 Multimodal Financial Benchmarks
Most financial benchmarks focus solely on text, with limited exploration of visual and audio data. Recent works like [56] address financial document parsing through OCR, while FinAudio [5] targets financial audio content such as earnings calls. However, a unified evaluation framework that integrates all three modalities remains absent. Our work introduces a benchmark that incorporates text, visual, and audio data, emphasizing two key directions: (1) dynamic dataset filtering to prioritize challenging content and (2) promoting multilingual and multimodal benchmarks to reflect real-world financial complexities.
For visual modality, existing financial datasets such as CORD [52] and FUNSD [29] focus on semi-structured documents but lack the complexity of financial statements. Applied systems like nvIngest [49] extract tabular data from PDFs but are not research benchmarks. Modern multimodal LLMs, such as LayoutLMv2 [83], effectively combine text, layout, and visual cues but focus mainly on entity extraction rather than complete document conversion. Our benchmark targets the end-to-end transformation of scanned documents into structured formats like HTML, aligning with practical financial analytics needs.
AudioLLMs such as AudioGPT [24] and AudioPaLM [60] excel in spoken content processing. Financial-specific benchmarks are scarce, with FinAudio [5] being the first to focus on financial audio but lacking integration with text and visual data. Our benchmark unifies text, visual, and audio modalities, enabling comprehensive evaluation of multimodal financial reasoning with metrics that assess both accuracy and semantic consistency across modalities.
Appendix B Structured Difficulty-Aware Benchmarking of English Datasets
Appendix C Limitations
@@
  • Incomplete Model Coverage. Due to computational and access constraints, we evaluate only a curated subset of representative models. While these span diverse modalities, access types, and language capabilities, broader coverage of emerging models—particularly open-access and region-specific LLMs—would further strengthen future iterations of the benchmark.
@@
Appendix D Potential Societal Impacts of the Work
D.1 Potential positive societal impacts
@@
D.2 Potential Negative Societal Impacts
@@
Appendix E Ethical Considerations and Usage Disclaimer
@@
Appendix F Textual Monolingual Benchmarks
F.1 English

### v3

@@ 3 Experimental Results
Figure 5: Performance across difficulty levels: Easy, Medium, Hard.
@@
Finding 1 for RQ1: Cross-modal and cross-lingual reasoning in 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: exposes clear limits of current LLMs. Building upon our structured, difficulty-aware benchmarking framework, 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: consists of 6 easy, 18 medium, and 12 hard datasets spanning three modalities and five economically critical languages across three linguistic settings. All multilingual, vision, and audio datasets fall into the medium or hard tiers, posing nontrivial challenges even for the most advanced models. The leading model, GPT-4o, achieves an overall score of only 46.01%, revealing clear limitations even at the frontier of multimodal and multilingual reasoning. The next best models, Qwen2.5-Omni and Llama-4, reach 33.53% and 20.89%, respectively, underscoring the steep difficulty gradient established by our benchmark. Notably, all three top-performing systems are multimodal and multilingual, reinforcing the necessity of these capabilities for robust financial reasoning. In contrast, monomodal and monolingual models perform considerably worse. The strongest text-only model, Llama-3.1-70B, attains merely 14.07% overall (with a best text-modality score of 42.20%). Likewise, modality- or language-specific models such as Whisper-V3 (audio-only, 17.19%), LLaVA-1.6 (vision-only, 5.15%), and FinMA-7B (English-only, 8.94%) trail far behind. These disparities highlight the inherent limitations of models lacking cross-modal and cross-lingual reasoning, emphasizing the critical need for integrated multimodal and multilingual understanding in financial AI, precisely the kind of comprehensive evaluation that 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is designed to provide.
Finding 2 for RQ2: Multimodal models exhibit trade-offs across modalities. Figure 6 illustrates the stratified performance across modalities. While GPT-4o leads overall performance on 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, it is surpassed by the text-specialized Llama-3.1-70B in text-only tasks (42.20% vs. 40.98%). Similarly, other leading multimodal models such as Qwen2.5-Omni and Llama-4 fall short of their text-only counterparts Qwen2.5-32B (38.07% vs. 33.06%) and Llama-3.1-70B (42.20% vs. 39.50%) in textual evaluations. This trade-off becomes more pronounced in mid-tier models like Gemma-3-4B and Gemma-3-27B, which show comparable vision-task performance (18.88% and 26.80%) to vision-only models (LLaVA-1.6: 15.45%, Deepseek-VL: 14.14%) but substantially lower text-task performance (12.74% and 15.74%). These trends highlight the difficulty of preserving text-specific optimization when expanding to multimodal capabilities. In contrast, multimodal models decisively outperform unimodal baselines in vision and audio tasks, i.e, GPT-4o achieves 46.02% in vision and 51.04% in audio, surpassing modality-specific counterparts. The asymmetric performance reveals that while text tasks benefit from data maturity, vision and audio gain more from multimodal integration, making unified models essential for complex financial real-world applications.
Finding 3 for RQ3: Scaling monolingual ability does not yield multilingual reasoning. Figure 6 illustrates the stratified performance across linguistic settings. Although GPT-4o consistently ranks among the top performers across individual monolingual settings, it performs poorly on multilingual tasks (9.79% on PolyFiQA-Easy and 5.31% on PolyFiQA-Expert; see detailed error analysis in Appendix I). A similar pattern holds for Qwen2.5-Omni (18.81% and 16.35%), suggesting that strong monolingual proficiency does not necessarily translate into multilingual reasoning. In contrast, models with lower overall benchmark performance, such as Deepseek-V3 (34.72% and 30.35%) and Llama-3.1-70B (25.04% and 18.56%), achieve markedly better results on multilingual tasks. These findings indicate that multilingual reasoning constitutes a fundamentally different challenge from monolingual evaluation, one that cannot be addressed simply by scaling monolingual capacity but instead requires dedicated modeling for multilingual understanding and reasoning. Given the increasing demand for simultaneous multilingual comprehension in global financial contexts, 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, particularly through its multilingual component (PolyFiQA-Easy and PolyFiQA-Expert, the first multilingual financial datasets), provides a critical stress test for this capability, capturing dimensions of multilingual reasoning that isolated monolingual evaluations fail to reveal.
Finding 4 for RQ4: Difficulty-aware benchmarking enables stratified, fine-grained, and dynamic evaluation. Figure 6 illustrates the stratified performance across difficulty levels. 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: introduces a three-tier structure (easy, medium, and hard) enabling layered evaluation of model capabilities. Performance converges at the hard tier, where all models struggle, whereas the easy and medium tiers yield clearer distinctions in capability. Except for Gemma-3-4B and Gemma-3-27B, which outperform on medium datasets due to their multimodal strengths, most models follow the monotonic trend easy>>medium>>hard. General-purpose models such as GPT-4o, Llama-4, and Qwen2.5-Omni demonstrate strong performance on easy and medium tiers but face substantial challenges on hard tasks. These hard-tier datasets (PolyFiQA-Easy, PolyFiQA-Expert, OCR-based, and other hard-tier datasets) require multilingual, multi-document reasoning and the joint interpretation of textual and structural information, posing a fundamental challenge for current models. By contrast, domain-specific models such as o3-mini, XuanYuan, and R1-Qwen32B-JA underperform on easy datasets, suggesting that heavy specialization may compromise generalization, a consideration for future training paradigms. Through this difficulty-aware design, 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: enables stratified and interpretable evaluation, revealing nuanced model strengths and weaknesses across complexity levels. As model capabilities advance, these tiers may shift dynamically, medium datasets may become easy, allowing the benchmark to evolve alongside model progress and real-world financial reasoning demands.
@@ 4 Conclusion
We present 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first multilingual and multimodal benchmark for evaluating real-world financial applications under difficulty-aware settings. Our results show persistent performance gaps across languages, modalities, and difficulty levels, with even GPT-4o exhibiting substantial limitations in complex scenarios. The newly introduced datasets (PolyFiQA and OCR) constitute the core contribution of 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: by addressing critical gaps in multilingual reasoning and document understanding from scanned financial reports. Together, they reveal weaknesses in cross-lingual and cross-modal reasoning. 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a comprehensive foundation for developing inclusive, multilingual, and multimodal financial LLMs capable of realistic expert-level reasoning.
Limitations
@@
  • Relatively Small Scale of Newly Introduced Datasets. While we introduce several novel datasets, their sample sizes are modest compared with large-scale benchmarks in general NLP. This constraint arises from the high annotation cost and domain expertise required for financial data. Future expansions with larger and more diverse samples will further enhance statistical robustness and task coverage.
  • Limited Language Coverage in the Audio Modality. Our audio evaluation currently focuses on English earnings calls. Due to regional, copyright, and legal restrictions, obtaining and releasing multilingual earnings-call recordings remains challenging. Nevertheless, the U.S. market is globally representative, and our dataset spans three years, multiple sectors, and diverse downstream tasks, providing a meaningful testbed for financial speech understanding. Future extensions will incorporate multilingual earnings-call data to broaden cross-lingual evaluation of audio-based financial reasoning.
  • Incomplete Model Coverage. Due to computational and access constraints, we evaluate only a curated subset of representative models. While these span diverse modalities, access types, and language capabilities, broader coverage of emerging models, particularly open-access and region-specific LLMs, would further strengthen future iterations of the benchmark.
@@
Potential Risks
Potential Positive Societal Impacts.
@@
Potential Negative Societal Impacts.
@@
Ethical Considerations and Usage Disclaimer
@@
Appendix A Related Work
A.1 Trend of Benchmark Development
With the rapid development of large language models, many traditional benchmark datasets have become insufficiently challenging to effectively distinguish between models of varying capacities Ivanov and Volkov (2025). Even relatively small-scale models now achieve comparable performance to much larger ones on these benchmarks, despite notable disparities in their real-world capabilities Hsieh et al. (2023). The FinBen benchmark Xie et al. (2024b) also reveal that certain datasets are unable to clearly distinguish differences in model capabilities by conducting thorough experiments and evaluates LLMs across diverse financial tasks.
This performance saturation has motivated increasing efforts to design more difficult and nuanced evaluation tasks across various domains He et al. (2024); Wang et al. (2024); White et al. (2024a). In the general domain, the BIG-Bench dataset Suzgun et al. (2022a) compiles challenging tasks to stress-test LLMs. Subsequent work examines whether prompting techniques can enhance performance in such tasks. In the medical domain, Chen et al. Chen et al. (2025a) developed QA tasks requiring clinical reasoning and explanation, providing a more rigorous evaluation of LLMs’ capabilities in healthcare. In the financial domain, Zhao et al. Zhao et al. (2023) introduced datasets of varying difficulty. For simpler tasks, small models perform comparably to larger ones. However, for complex benchmarks like Complong, involving multi-table reasoning and long document processing, smaller models struggle while larger models maintain a lead. This highlights the need for more discriminative benchmarks to expose nuanced differences in model capabilities.
Building on these observations and the proliferation of financial datasets, some recent work has proposed criteria to filter and prioritize representative datasets for benchmarking Reuel et al. (2024); Cao et al. (2025a). These efforts assess benchmarks based on implementation, design, documentation, maintenance, and construction to determine their effectiveness. However, these approaches do not systematically evaluate the impact of dataset difficulty or establish objective criteria for filtering datasets based on complexity.
A.2 Multilingual Textual Financial Benchmarks
Despite the growing number of financial benchmarks developed to assess the capabilities of large language models in domain-specific tasks, most existing benchmarks remain monolingual, predominantly focusing on English or Chinese. Notable examples include PIXIU Xie et al. (2023), CFBenchmark Lei et al. (2023), FinanceBench Islam et al. (2023), and BizBench Koncel-Kedziorski et al. (2023), each offering tasks such as question answering, text classification, and numerical reasoning, all rooted in financial documents or news within a single language setting. Some efforts have extended financial benchmarking to additional languages. For instance, Peng et al. (2025) and Hirano (2024) introduced benchmarks that encompass Greek and Japanese, broadening the linguistic focus beyond English and Chinese. Although works like Xie et al. (2024b) and Zhang et al. (2024b) expand beyond English by including tasks in both English and Spanish, they treat the two languages as separate tasks rather than integrating them into a unified multilingual evaluation. This approach limits the ability to assess cross-lingual generalization within financial contexts.
Some recent studies have attempted to expand financial benchmarks by incorporating bilingual tasks. For example, DolFin Nakhlé et al. (2025) includes English-Chinese financial QA and reasoning tasks. However, its focus remains limited to translation-based tasks, which do not fully reflect real-world multilingual applications.
Currently, there is a lack of comprehensive financial benchmarks that simultaneously encompass monolingual, bilingual, and multilingual settings. Most existing resources either focus solely on monolingual tasks or limited bilingual tasks, without integrating them into a unified multilingual framework. This gap highlights the need for more inclusive and versatile evaluation resources that can assess model performance across diverse linguistic settings. Our work addresses this limitation by proposing a benchmark that not only evaluates LLMs across a wider range of financial tasks, but also considers dataset value filtering and cross-lingual robustness, setting the stage for more comprehensive and discriminative benchmarking in the financial domain.
A.3 Multimodal Financial Benchmarks
Most financial benchmarks focus solely on text, with limited exploration of visual and audio data. Recent works like Qiu et al. (2024) address financial document parsing through OCR, while FinAudio Cao et al. (2025b) targets financial audio content such as earnings calls. However, a unified evaluation framework that integrates all three modalities remains absent. Our work introduces a benchmark that incorporates text, visual, and audio data, emphasizing two key directions: (1) dynamic dataset filtering to prioritize challenging content and (2) promoting multilingual and multimodal benchmarks to reflect real-world financial complexities.
For visual modality, existing financial datasets such as CORD Park et al. (2019) and FUNSD Jaume et al. (2019) focus on semi-structured documents but lack the complexity of financial statements, FinMME Luo et al. (2025) and MME-Finance Gan et al. (2024) focus on image understanding rather than the complex structure of document-based financial information. Applied systems like nvIngest NVIDIA Research (2024) extract tabular data from PDFs but are not research benchmarks. Modern multimodal LLMs, such as LayoutLMv2 Xu et al. (2021), effectively combine text, layout, and visual cues but focus mainly on entity extraction rather than complete document conversion. Our benchmark targets the end-to-end transformation of scanned documents into structured formats like HTML, aligning with practical financial analytics needs.
AudioLLMs such as AudioGPT Huang et al. (2024b) and AudioPaLM Rubenstein et al. (2023) excel in spoken content processing. Financial-specific benchmarks are scarce, with FinAudio Cao et al. (2025b) being the first to focus on financial audio but lacking integration with text and visual data. Our benchmark unifies text, visual, and audio modalities, enabling comprehensive evaluation of multimodal financial reasoning with metrics that assess both accuracy and semantic consistency across modalities.
Appendix B Task Categories
Inspired by FinBen Xie et al. (2024a), we organize candidate datasets under a unified taxonomy of seven core financial NLP tasks:
  • Information Extraction (IE) focuses on converting unstructured financial text into structured outputs.
  • Textual Analysis (TA) assesses a model’s ability to interpret sentiment, topic, or tone in financial discourse.
  • Question Answering (QA) evaluates comprehension of financial content through question answering.
  • Text Generation (TG) focuses on producing coherent, informative, and factually accurate financial text.
  • Risk Management (RM) targets detection or analysis of risk-related signals.
  • Forecasting (FO) measures a model’s ability to predict market trends or investor behavior.
  • Decision-Making (DM) simulates complex financial decision processes.
Appendix C Evaluation Models
Access Modality Language Domain Target Language Model MOF Class Closed Multimodal Multilingual General English GPT-4o Hurst et al. (2024) Class III - Open Model - In progress (17%) Text Multilingual General English o3-mini OpenAI (2025) Class III - Open Model - In progress (0%) Open Multimodal Multilingual General English meta-llama/Llama-4-Scout-17B-16E-Instruct AI (2025) Class III - Open Model - In progress (0%) General English google/gemma-3-4b-it Team (2025) Class III - Open Model - In progress (17%) General English google/gemma-3-27b-it Team (2025) Class III - Open Model - In progress (17%) General English Qwen/Qwen2.5-Omni-7B Xu et al. (2025) Class III - Open Model - In progress (67%) Text Multilingual General English meta-llama/Llama-3.1-70B-Instruct Dubey et al. (2024) Class III - Open Model - In progress (0%) General English Deepseek-V3 Liu et al. (2024) Class III - Open Model - In progress (17%) General English Qwen/Qwen2.5-32B-Instruct Yang et al. (2025) Class III - Open Model - In progress (33%) General Japanese cyberagent/DeepSeek-R1-Distill-Qwen-32B-Japanese Ishigami (2025) Class III - Open Model - In progress (50%)1 Financial Chinese Duxiaoman-DI/Llama3.1-XuanYuan-FinX1-Preview Team (2024) Class III - Open Model - In progress (0%)1 Financial Spanish TheFinAI/FinMA-ES-Bilingual Zhang et al. (2024a) Class III - Open Model - In progress (50%)1 Financial Greek TheFinAI/plutus-8B-instruct Peng et al. (2025) Class III - Open Model - In progress (17%)1 English Financial English TheFinAI/finma-7b-full Xie et al. (2023) Class III - Open Model - In progress (67%)1 Vision Multilingual General English Deepseek-VL-7B-Chat Lu et al. (2024) Class III - Open Model - In progress (17%) English General English llava-hf/llava-v1.6-vicuna-13b-hf Liu et al. (2023a) Class III - Open Model - Qualified & Class II - Open Tooling - In progress (80%)1 Audio Multilingual General English Whisper-V3 Radford et al. (2023) Class III - Open Model - In progress (67%) General English Qwen2-Audio-7B Chu et al. (2024) Class III - Open Model - In progress (67%) General English Qwen2-Audio-7B-Instruct Chu et al. (2024) Class III - Open Model - In progress (67%) General English SALMONN-7B Tang et al. (2024) Class III - Open Model - In progress (50%)1 General English SALMONN-13B Tang et al. (2024) Class III - Open Model - In progress (50%)1
  • 1
Appendix D Structured Difficulty-Aware Benchmarking of English Datasets
Appendix E Textual Monolingual Benchmarks
E.1 English